OpenStudy (anonymous):

An equilateral triangle and a rectangle have the same perimeter. The width of the rectangle is three inches longer than the sides of the triangle. If the rectangle is eight inches tall, what is the area of the rectangle?

OpenStudy (anonymous):

translates to 2a+2b=3c a=c+3 b=8